### v2

# Maximum principles for matrix-valued regular functions of a quaternionic variable

Abstract.
A quaternionic matrix-valued regular function is a map $F:\Omega\rightarrow M_{n}(\mathbb{H})$ whose entries are regular functions of a quaternion variable, where $\Omega$ is a domain in $\mathbb{H}$. Our aim is to bring out some maximum norm principles for such functions. We derive a decomposition theorem for such functions and also prove a Caratheodory-Rudin type approximation theorem for functions in the quaternionic Schur class. This in turn yields that a $2\times 2$ norm one matrix-valued function can be approximated by quaternionic rational inner functions.
Key words and phrases:
Quaternions; symmetric slice domain; matrix-valued regular functions; complex adjoint matrix; maximizing vector; maximum principle; rational inner functions
2020 Mathematics Subject Classification:
Primary: 30G35, 15B33, 47S05 Secondary: 15A18
Corresponding author
1. A brief introduction
We work throughout over the fields $\mathbb{R}$ and $\mathbb{C}$ of real and complex numbers and the skew field / division ring of the set of real quaternions $\mathbb{H}$, defined by, $\mathbb{H}:=\{q=a_{0}+a_{1}i+a_{2}j+a_{3}k:i^{2}=j^{2}=k^{2}=ijk=-1,ij=-ji=k,jk=-kj=i,ki=-ik=j,\ a_{i}\in\mathbb{R}\}$. For $q=a_{0}+a_{1}i+a_{2}j+a_{3}k\in\mathbb{H},\ |q|:=\sqrt{a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}}$ denotes the modulus of $q$ and $\overline{q}:=a_{0}-a_{1}i-a_{2}j-a_{3}k$ denotes the conjugate of $q$. Quaternions $p$ and $q$ are similar if $s^{-1}qs=p$ for some $s\in\mathbb{H}\setminus\{0\}$. If quaternions $p$ and $q$ are similar, then $|p|=|q|$. Let $\mathbb{H}^{n}$ denote the set of all column vectors with entries from the quaternions. For $v=\begin{bmatrix}v_{1}\\ \vdots\\ v_{n}\end{bmatrix}\in\mathbb{H}^{n}$, $v_{j}\in\mathbb{H}$ and $\alpha\in\mathbb{H}$, let $\alpha v=\begin{bmatrix}\alpha v_{1}\\ \vdots\\ \alpha v_{n}\end{bmatrix}$. With respect to the standard addition and the scalar multiplication defined above, $\mathbb{H}^{n}$ is a left quaternion vector space over the division ring $\mathbb{H}$. Standard linear algebra concepts such as linear independence, spanning sets, basis, direct sums, and so on work in exactly the same way as for finite dimensional vector spaces over commutative division rings (see [12] for details). The subspace spanned by $v_{1},v_{2},\ldots,v_{n}\in\mathbb{H}^{n}$ is denoted by $\text{span}_{\mathbb{H}}\{v_{1},v_{2},\ldots,v_{n}\}=\{v_{1}q_{1}+v_{2}q_{2}+\cdots+v_{n}q_{n}:q_{i}\in\mathbb{H}\}$. The inner product on $\mathbb{H}^{n}$ is defined by $\langle u,v\rangle=v^{*}u$ for $u,v\in\mathbb{H}^{n}$; that is, for any two vectors $u,v\in\mathbb{H}^{n}$ and $a\in\mathbb{H}$, we have $\langle u,v.a\rangle=\langle u,v\rangle.a$ and $\langle u.a,v\rangle=\bar{a}.\langle u,v\rangle$. Thus, in particular, $\mathbb{H}$ is a topological space with the Euclidean topology , where the norm of an element $q$ is given by $\|q\|^{2}=x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$.
We shall denote by $M_{n}(\mathbb{C})$ and $M_{n}(\mathbb{H})$ the rings of $n\times n$ matrices with complex and quaternion entries respectively. Recall that any vector $x=y_{0}+y_{1}i+y_{2}j+y_{3}k\in\mathbb{H}^{n}$, where $y_{0},y_{1},y_{2},\ \&\ y_{3}\in\mathbb{R}^{n}$, can be uniquely expressed as $x=(y_{0}+y_{1}i)+(y_{2}+y_{3}i)j\ =\ x_{1}+x_{2}j$, where $x_{1}$ and $x_{2}\in\mathbb{C}^{n}$. Similarly, any quaternion matrix $A=B_{0}+B_{1}i+B_{2}j+B_{3}k\in M_{n}(\mathbb{H})$, where $B_{0},B_{1},B_{2},\ \&\ B_{3}\in M_{n}(\mathbb{R})$, can be uniquely expressed as $A=(B_{0}+B_{1}i)+(B_{2}+B_{3}i)j=A_{1}+A_{2}j$, where $A_{1}$ and $A_{2}\in M_{n}(\mathbb{C})$. This suggests that one can associate a complex block matrix to a given quaternion matrix as given below.
Definition 1.1.
For $A\in M_{n}(\mathbb{H})$ that is expressed as $A=A_{1}+A_{2}j$ for some $A_{1},A_{2}\in M_{n}(\mathbb{C})$, the $2n\times 2n$ block matrix $\begin{bmatrix}A_{1}&A_{2}\\ -\overline{A_{2}}&\overline{A_{1}}\end{bmatrix}$ is called the complex adjoint matrix of $A$ and is denoted by $\mbox{\Large$\chi$}_{A}$.
It can be easily verified that $A\in M_{n}(\mathbb{H})$ is invertible if and only if its complex adjoint matrix $\mbox{\Large$\chi$}_{A}$ is invertible. The Frobenius norm and the operator norm of a quaternion matrix are defined by
[…]
respectively.
[…]
2. Definitions and Preliminaries
Necessary definitions, facts, terminology and theorems used further in the paper appear in this section. We begin by stating that throughout the paper, a domain in $\mathbb{H}$ is an open connected set with respect to Euclidean topology. The unit ball in $\mathbb{H}$, centered at the origin will be denoted by $\mathbb{B}$.
Definition 2.1.
The $2$-sphere of quaternionic imaginary units, denoted by, $\mathbb{S}$, is defined as
$\mathbb{S}=\{q\in\mathbb{H}:q^{2}=-1\}=\{x_{1}i+x_{2}j+x_{3}k:x_{1}^{2}+x_{2}^{2}+x_{3}^{2}=1\}.$
We shall use the above notation $\mathbb{S}$ throughout this paper to denote only the $2$-sphere.
For $I\in\mathbb{S}$, we denote by $L_{I}$ the set $\mathbb{R}+\mathbb{R}I$, which we may identify with $\mathbb{C}$. We then define $\Omega_{I}$ to be the intersection of any domain $\Omega$ with $L_{I}$ and $f|_{I}$ to be the restriction of $f:\Omega\rightarrow\mathbb{H}$ on $\Omega_{I}$. A function $f:\Omega\rightarrow\mathbb{H}$ is said to be regular on a domain $\Omega$ if for any $I\in\mathbb{S}$, the function $f|_{I}$ is complex holomorphic on $\Omega_{I}$. Every regular function on $\mathbb{B}$ admits a nice power series representation, as defined below.
Definition 2.2.
A function $f:\mathbb{B}\rightarrow\mathbb{H}$ is said to be a left regular function if it admits a power series representation as $f(q)=\displaystyle\sum_{n=0}^{\infty}q^{n}f_{n}$, where $f_{n}\in\mathbb{H}$ and $\overline{\displaystyle{\lim_{n\to\infty}}}\sqrt[n]{|f_{n}|}\leq 1$.
An analagous definition of right regular functions exist; however, we work only with left regular functions in this paper. All of the definitions, facts and theorems stated below can be found in [8] or [9]. With the above definitions, any left polynomial over quaternions is an example of a regular function. Another simple example of a regular function over the quaternions is $\frac{1}{q}$, where $0\neq q\in\mathbb{H}$. There are functions that are not regular too. We shall exhibit one in Example 3.13.
Remark: From now on, we use the phrase regular function only to mean that it is a left regular function, to avoid confusion, although the definitions and facts presented below also hold for right regular functions.
We now recall the Splitting Lemma (see for instance [8] or [9]), which helps us to restrict a regular function to $L_{I}$, and hence get complex holomorphic function.
Lemma 2.3 (Splitting lemma).
Let $f$ be a regular function defined on a domain $\Omega$. Then for any $I\in\mathbb{S}$ and any $J\in\mathbb{S}$ with $I\perp J$, there exists two holomorphic functions $F,G:\Omega_{I}\rightarrow L_{I}$ such that for every $z=x+yI$,
$f_{I}(z)=f|_{\Omega_{I}}(z)=F(z)+G(z)J.$
We now proceed to the definition of a slice domain in $\mathbb{H}$.
Definition 2.4.
Let $\mathbb{S}$ be the $2$-sphere from Definition 2.1 and let $\Omega$ be a domain in $\mathbb{H}$ that intersects the real axis. $\Omega$ is called a slice domain if for all $I\in\mathbb{S}$, the intersection $\Omega_{I}=\Omega\cap L_{I}$ is a domain in $L_{I}$.
Slice domains play an important role in proving results such as the identity principle as well as in ensuring continuity of regular functions in a domain.
Definition 2.5.
A subset $T$ of $\mathbb{H}$ is called axially symmetric if for all points $x+yI\in T$ with $x,y\in\mathbb{R}$ and $I\in\mathbb{S}$ (the $2$-sphere), the set $T$ contains the whole sphere $x+y\mathbb{S}$.
The unit open ball $\mathbb{B}$ is an example of a symmetric slice domain in $\mathbb{H}$. Therefore, some of the results we prove for a symmetric slice domain will also be true when the domain is the unit ball $\mathbb{B}$. We now state below the identity principle for regular functions on a slice domain $\Omega$ (see Theorem $1.12$ of [9]).
Theorem 2.6 (Identity principle).
Let $f$ and $g$ be regular functions on a slice domain $\Omega$. If $f$ coincides with $g$ on a subset of the intersection $\Omega_{I}=L_{I}\cap\Omega$ having an accumulation point in $\Omega_{I}$ for some $I$ in the $2$-sphere $\mathbb{S}$, then $f\equiv g$ in $\Omega$.
The following definition is that of a regular matrix-valued function.
Definition 2.7.
Let $F:\Omega\rightarrow M_{n}(\mathbb{H})$ be a quaternionic matrix-valued function defined on a domain $\Omega$ in $\mathbb{H}$. $F$ is said to be regular if each $ij$-entry of $F$ is a regular function on $\Omega$.
For example, a left matrix polynomial, which is defined on the whole of $\mathbb{H}$, is a simple example of a regular matrix-valued function. In a recent work, Condori [7] obtained interesting results concerning extensions of the maximum modulus theorem for matrix-valued analytic functions on a domain in $\mathbb{C}$. It turns out that the maximum modulus principle holds for regular functions of a quaternion variable over a slice domain (see for instance Chapter $7$ of [9]). The purpose of this paper is to investigate to what extent the maximum principle holds for quaternionic matrix-valued regular functions on a domain $\Omega$ in $\mathbb{H}$.
We now proceed to give the definitions of a few more standard notions that will be used in the later sections. The first of these is the definition of regular extension of $f_{I}$.
Definition 2.8.
Let $\Omega$ be a symmetric slice domain and let $I\in\mathbb{S}$. If $f_{I}:\Omega_{I}\rightarrow\mathbb{H}$ is holomorphic, then there exists a unique regular function $g:\Omega\rightarrow\mathbb{H}$ such that $g_{I}=f_{I}$ in $\Omega_{I}$. Such a function $g$ will be denoted by $ext(f_{I})$ and is called the regular extension of $f_{I}$ and given by
$g(q)=ext(f_{I}(x+yJ))=\frac{1}{2}\Big[f(x+yJ)+f(x-yJ)+IJ[f(x-yJ)-f(x+yJ)]\Big].$
The following three definitions (in that order) are those of the regular product of two regular functions, regular conjugate and symmetrization of a function and the regular reciprocal of a regular function, respectively.
Definition 2.9.
Let $f,g$ be regular functions on a symmetric slice domain $\Omega$. For some $I,J\in\mathbb{S}$ with $I\perp J$, let $F,G,H,K$ be holomorphic functions from $\Omega_{I}$ to $L_{I}$ such that $f_{I}=F+GJ,\ \&\ g_{I}=H+KJ$. Consider the holomorphic function defined on $\Omega_{I}$ by The regular product of $f$ and $g$, denoted by fgf*g, is defined to be $f*g=ext(f_{I}*g_{I})$, where
$f_{I}*g_{I}(z)=[F(z)H(z)-G(z)\overline{K(\bar{z})}]+[F(z)K(z)+G(z)\overline{H(\bar{z})}]J.$
Definition 2.10.
Let $f:\mathbb{B}(0,R)\rightarrow\mathbb{H}$ be a regular function defined on a ball of radius $R$ and let $f(q)=\displaystyle\sum_{n\in\mathbb{N}}q^{n}a_{n}$ be its power series expansion. The regular conjugate of $f$ is the regular function defined by
$f^{c}(q):=\displaystyle\sum_{n\in\mathbb{N}}q^{n}\overline{a_{n}}$
on the same ball $\mathbb{B}(0,R)$. The symmetrization of $f$ is the function
$f^{s}=f*f^{c}=f^{c}*f.$
Remark 2.11.
For symmetric slice domains, the above definition can be suitably modified, where the conjugate is defined as in Definition $1.3.4$ of [9].
Definition 2.12.
Let $f:\Omega\rightarrow\mathbb{H}$ be a regular function. Then the regular reciprocal is the function $f^{-*}:=\frac{1}{f^{s}}f^{c}$.
The final definition in this section is that of the $\sigma$-ball in $\mathbb{H}$.
Definition 2.13.
For all $p,q\in\mathbb{H}$, let
[…]
The $\Sigma$-ball of radius $R>0$ is the set
$\Sigma(p,R)=\{q\in\mathbb{H}:\sigma(q,p)<R\}.$
We state the following fact before describing the results obtained in this paper.
Fact 2.14.
As pointed out in Theorem $2.12$ of [9], if $\Omega$ is an arbitrary domain in $\mathbb{H}$, then any regular function will have a power series expansion in each $\Sigma$-ball in $\Omega$.
A brief description of the results obtained are as below. After a brief introduction, the necessary definitions and preliminaries are brought out. This is followed by the section on main results, which is subdivided into four subsections for ease of reading. The first of these concerns maximum principles for Frobenius and operator norms of matrix-valued functions. Maximizing vectors and a decomposition result for matrix-valued regular functions are brought out next. We then discuss maximum principles involving singular values of matrix-valued functions. The last subsection concerns approximating a regular matrix-valued function by rational inner functions. We bring out a Caratheodory-Rudin type approximation result of a regular function of a quaternion variable and use it to prove that a $2\times 2$ matrix-valued regular function of a quaternion variable can be approximated by matrix-valued rational inner functions.
3. Main Results
The main results are presented in this section. This section is subdivided into subsections, for ease of reading.
3.1. The Frobenius and Operator norm case
We begin by proving - very similar to the complex case (see Theorem $1$, [7]) - that the maximum modulus principle holds for quaternionic matrix-valued regular functions when working with the Frobenius norm.
Theorem 3.1.
Let $\Omega$ be a domain in $\mathbb{H}$ and $F:\Omega\rightarrow M_{n}(\mathbb{H})$ be a quaternionic matrix-valued regular function. Suppose there exists a point $q_{0}\in\Omega_{I}(=\Omega\cap L_{I})$ for some $I\in\mathbb{S}$ such that $\|F(q)\|_{F}\leq F(q_{0})$ for every $q\in\Omega$. Then $F(q)$ is a constant.
Proof.
Let the complex adjoint matrix of $F(q)=G(q)+H(q)j$ be denoted by $\widetilde{F}(q)$. The Frobenius norm of $F(q)$ is,
$\|F(q)\|_{F}^{2}=\text{trace}(F(q)F(q)^{*})=|f_{11}|^{2}+|f_{12}|^{2}+|f_{21}|^{2}+|f_{22}|^{2}.$
We now calculate the Frobenius norm of the complex adjoint matrix $\widetilde{F}(q)$. Then we get
$\|\widetilde{F}(q)\|_{F}^{2}=2(|g_{11}|^{2}+|g_{12}|^{2}+|g_{21}|^{2}+|g_{22}|^{2}+|h_{11}|^{2}+|h_{12}|^{2}+|h_{21}|^{2}+|h_{22}|^{2}).$
We also have that $f_{mn}=g_{mn}+h_{mn}j$ so that $|f_{mn}|^{2}=|g_{mn}|^{2}+|h_{mn}|^{2}$. Therefore, $\|\widetilde{F}(q)\|_{F}^{2}=2(\|F(q)\|_{F}^{2})$ and hence $\|F(q)\|_{F}\leq F(q_{0})$, thereby implying that $\|\widetilde{F}(q)\|_{F}\leq\sqrt{2}F(q_{0})$. Then, by the maximum Frobenius norm principle for matrix-valued analytic functions, as proved in Theorem $1$ of [7], it follows that $\widetilde{F}(q)$ is constant and therefore $F(q)$ is also a constant. ∎
The above result need not be true in general for any matrix norm. The example given in [7] works in the quaternion case too. Before proceeding further we state the Hahn-Banach theorem and one of its consequences proved by Suhomlinov for the quaternion case (see [13] or Theorem $4.10.1$ & Corollary $4.10.2$ of [6] for details).
Theorem 3.2.
Let $L\subset\mathbb{H}$ be a left linear space over $\mathbb{H}$ and $M$ a subspace of $L$. Suppose $f:M\rightarrow\mathbb{H}$ is a left linear functional whose domain is $M$ and has the property $|f(q)|\leq p(q)$ for all $q\in M$, where $p:L\rightarrow\mathbb{R}$ is a real valued functional defined on all of $L$ with the properties:
  • (1)
    $p(q)\geq 0$,
  • (2)
    $p(q+r)\leq p(q)+p(r)$, and
  • (3)
    $p(\alpha q)=p(q)|\alpha|$.
[…]
We now use the above theorem to prove our next result, which is again along the same lines as that of Theorem $2$ from [7].
Theorem 3.3.
Let $\mathbb{B}$ be the unit ball centered at the origin in $\mathbb{H}$ and $F:\mathbb{B}\rightarrow M_{n}(\mathbb{H})$ be a quaternionic matrix-valued regular function. If $\|F(q)\|$ attains its maximum in $\mathbb{B}$, then $\|F(q)\|$ is a constant on $\mathbb{B}$.
Proof.
Let $q_{0}\in\mathbb{B}$ be such that $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\mathbb{B}$. Let $\|F(q_{0})\|\neq 0$. From Theorem 3.2, it follows that there is a bounded linear functional $T:M_{n}(\mathbb{H})\rightarrow\mathbb{H}$ such that
$T(F(q_{0}))=\|F(q_{0})\|\text{ and }\|T\|=1.$
Since $F$ is left regular on $\mathbb{B}$, it admits a power series expansion as
$F(q)=\sum_{n=0}^{\infty}q^{n}F_{n}.$
$T$ being a bounded (left) linear transformation, we have
$T(F(q))=T(\sum_{n=0}^{\infty}q^{n}F_{n})=\sum_{n=0}^{\infty}q^{n}T(F_{n}).$
Thus, $T(F):\mathbb{B}\rightarrow\mathbb{H}$ is regular on $\mathbb{B}$, and
$|T(F(q))|$ $\leq\|F(q)\|$
$\leq\|F(q_{0})\|$
$=T(F(q_{0})).$
It now follows from the maximum principle for regular functions of a quaternion variable (see Theorem $7.1$ of [9]) that $T(F(q))$ is constant on $\mathbb{B}$. Then,
$\|F(q)\|\geq|T(F(q))|=|T(F(q_{0}))|=\|F(q_{0})\|\geq\|F(q)\|\ \text{for all}\ \ q\in\mathbb{B}.$
Hence $\|F(q)\|$ is a constant on $\mathbb{B}$. ∎
Remark 3.4.
As pointed out in in Theorem $2.12$ of [9], if we work with an arbitrary domain $\Omega$ in $\mathbb{H}$, then any regular function will have a power series expansion in each $\Sigma$-ball in $\Omega$ and in such a case, the proof of Theorem 3.3 can be modified appropriately.
3.2. Maximizing vectors and a decomposition theorem
We now define the notion of a maximizing vector over quaternions and prove a maximum operator norm principle for a quaternionic matrix-valued function. Throughout this section, the notation $\|\cdot\|$ will denote the operator norm of the matrix-valued function.
Definition 3.5.
A unit vector $q_{0}\in\mathbb{H}^{n}$ is called a maximizing vector for $A\in M_{n}(\mathbb{H})$ if
$\|Aq_{0}\|_{2}=\|A\|.$
$\mathbb{H}$ being a Euclidean space, the unit ball in $\mathbb{H}$ is a compact set. This assures that the maximizing vector always exists. We now prove that if $F(q)$ is a non constant matrix valued regular function whose operator norm attains its maximum at some point in $\Omega_{I}$, then there exists a vector $x_{0}$ such that $F(q)x_{0}$ is a constant.
Theorem 3.6.
Let $\Omega$ be a slice domain in $\mathbb{H}$ and $F:\Omega\rightarrow M_{n}(\mathbb{H})$ a regular matrix-valued function. If there exists a $q_{0}$ in $\Omega_{I}$ such that $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\Omega$ and $x_{0}$ is the maximizing vector for $F(q_{0})$, then $F^{k}(q_{0})x_{0}=0$ for all $k\geq 1$.
Proof.
Let the assumptions stated above hold. Choose $R>0$ such that $\Sigma(q_{0},R)$ contained in $\Omega$ and
$F(q)=\displaystyle\sum_{n\in\mathbb{N}}(q-q_{0})^{*n}F_{n},$
where $F_{n}\in M_{n}(\mathbb{H})$. Let $x\in\mathbb{H}$. Since $\mathbb{H}$ is a left Hilbert space, we have
[…]
$=\left\langle\displaystyle\sum_{n\in\mathbb{N}}(q-q_{0})^{*n}F_{n}x,\displaystyle\sum_{k\in\mathbb{N}}(q-q_{0})^{*k}F_{k}x\right\rangle$
$=\displaystyle\sum_{n\in\mathbb{N}}\displaystyle\sum_{k\in\mathbb{N}}(q-q_{0})^{*n}\langle F_{n}x,F_{k}x\rangle\overline{(q-q_{0})^{*k}}.$
Let $\delta_{I}(q_{0},R)\subseteq\Sigma(q_{0},R)$ be the disk in $L_{I}$ of radius $R$ and centered at $q_{0}$. Then for any $q\in\delta_{I}(q_{0},R)$, choose $r<R$ such that $q=q_{0}+re^{it}$. We then have,
[…]
By the regular power series representation (see [9]), $\|F(q)\|\leq\|F(q_{0})\|=\|F_{0}\|$ for all $q\in\delta_{I}(q_{0},R)$; in particular, for the vector $x$ and $r<R$, it follows that
[…]
Replacing $x$ by the maximizing vector $x_{0}$, we get
[…]
This implies that $F_{k}x_{0}=0$ for all $k\geq 1$ and hence $F^{k}(z_{0})x_{0}=0$ for all $k\geq 1$ and hence $F(q)x_{0}=F_{0}x_{0}=F(q_{0})x_{0}$ for all $q\in\delta_{I}(q_{0},R)$. We finally conclude from the identity principle (Theorem 2.6) that $F(q)x_{0}$ is constant on $\Omega$, as required. ∎
We now proceed to prove a decomposition for matrix-valued regular functions that is very similar to the singular value decomposition for complex matrices. We restrict our domain to be a symmetric slice domain in $\mathbb{H}$. We begin with a lemma, whose proof is routine.
Lemma 3.7.
Let $X$ and $Y$ be two $n\times n$ unitary matrices in $\mathbb{H}$ then for any matrix $A\in M_{n}(\mathbb{H})$,
$\|XAY\|=\|A\|.$
We now prove the aforesaid decomposition theorem.
Theorem 3.8.
Let $\Omega$ be a symmetric slice domain in $\mathbb{H}$ and let $F:\Omega\rightarrow M_{n}(\mathbb{H})$ be regular. If $q_{0}\in\Omega_{I}$ is such that $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\Omega$, then there are $n\times n$ constant unitary matrices $U$ and $V$ and a regular matrix valued function $G:\Omega\rightarrow M_{n-1}(\mathbb{H})$ such that
$F(q)=U\begin{bmatrix}\|F(q_{0})\|&0\\ 0&G(q)\end{bmatrix}V.$
Proof.
Assume that $\|F(q_{0})\|=1$ and let $x_{0}$ be the maximizing vector of $F(q_{0})$. Since $\|F(q)\|\leq\|F(q_{0})\|$, by Theorem 3.6 $F(q_{0})x_{0}$ is a constant on $\Omega$ and $F(q)x_{0}=F(q_{0})x_{0}$ for all $q\in\Omega$. Let us define a vector $y_{0}:=F(q_{0})x_{0}$ such that
$\|y_{0}\|^{2}=y_{0}^{*}y_{0}=x_{0}^{*}F(q_{0})^{*}F(q_{0})x_{0}=1.$
Observe that $F(q)x_{0}=F(q_{0})x_{0}$ implies that $y_{0}^{*}F(q)x_{0}=y_{0}^{*}F(q_{0})x_{0}$ for all $q\in\Omega$. Let $X_{0}$ and $Y_{0}$ be two unitary matrices whose first columns are $x_{0}$ and $y_{0}$ respectively. ($\mathbb{H}$ being a left Hilbert space, it is always possible to construct a unitary matrix from a norm one vector). Therefore we get,
$Y_{0}^{*}F(q)X_{0}=\begin{bmatrix}f_{11}(q)&f_{12}(q)\\ f_{21}(q)&f_{22}(q)\end{bmatrix}$
such that $f_{11}(q)=y_{0}^{*}F(q_{0})x_{0}=1$. Since $X_{0}$ and $Y_{0}$ are unitary matrices,
$\|Y_{0}^{*}F(q)X_{0}\|=\|F(q)\|=1.$
The operator norm of an $n\times n$ matrix being an upper bound on the euclidean norm of its columns and rows, we have $f_{12}(q)=f_{21}(q)=0$. We thus conclude that
$F(q)=Y_{0}\begin{bmatrix}1&0\\ 0&f_{22}(q)\end{bmatrix}X_{0}^{*},$
where $f_{22}(q)$ is a regular $(n-1)\times(n-1)$ matrix-valued function. We set $U=Y_{0},V=X_{0}^{*}$ and $G(q)=f_{22}(q)$ to get the desired conclusion. ∎
A few remarks are in order.
Remark 3.9.
  • Note that for $n=1$, the function $F:\Omega\rightarrow M_{1}(\mathbb{H})\cong\mathbb{H}$ is a regular function. Suppose there exists $q_{0}\in\Omega_{I}$ such that $\|F(q)\|\leq\|F(q_{0})\|$, for all $q\in\Omega$, then by maximum modulus principle (Theorem $7.1$ of [9]), it follows that $F$ is a constant on $\Omega$. By Theorem 3.8 we can write $F(q)$ as
    $F(q)=u\|F(q_{0})\|v$
    where u=v=1||u||=||v||=1. Let $x_{0}$ be the maximizing vector of $F(q_{0})$ so that
    $\|F(q_{0})\|=\sup_{\|x\|_{2}=1}\|F(q_{0})x\|_{2}=\|F(q_{0})x_{0}\|_{2}.$
    Then $u=x_{0},v=F(q_{0})x_{0}$, so that $F(q)=x_{0}\|F(q_{0})\|F(q_{0})x_{0}$.
  • Let $n=2$ and $F:\Omega\rightarrow M_{2}(\mathbb{H})$ a regular function. Suppose there exists $q_{0}\in\Omega_{I}$ such that $\|F(q)\|\leq\|F(q_{0})\|$ so that a decomposition as stated in Theorem 3.8 holds. If there exists a sequence of non-vanishing regular functions $g_{n}:\Omega\to\mathbb{H}$ which converges uniformly to $g:\Omega\to\mathbb{H}$, then by Hurwitz theorem for regular functions, $g$ is either identically zero or non-vanishing (see Theorem $4.5$ of [10]). In the former case, the function $F$ becomes a constant. Notice that there is a sequence of polynomials on $\Omega$ that converges uniformly to a given regular function on $\Omega$ (see for instance Chapter $3$ of [8]).
The following example shows that above decomposition holds only if such a vector $q_{0}$ exists in the domain $\Omega$.
Example 3.10.
Let $r$ and $R$ be positive reals numbers with $r<R$ and let $\Omega$ be the annular region defined by $\Omega:=\{q\in\mathbb{H}:r<|q|<R\}$. Consider the quaternionic matrix-valued function $F$ defined on $\Omega$ by $F(q)=\begin{bmatrix}q-1&2\\ 0&\dfrac{1}{q}\end{bmatrix}$. Since $q\neq 0$, $F$ is regular. By calculating the norm of $F(q)$, we get
$\|F(q)\|=\frac{\sqrt{|q-1|^{2}+4+\frac{1}{|q|^{2}}+\sqrt{(|q-1|^{2}+4+\frac{1}{|q|^{2}})^{2}-4|q-1|^{2}}}}{2}.$
It is then clear that $\|F(q)\|$ is non-constant, whose maximum is obtained at a point $q_{0}$ such that $|q_{0}|=r$. Suppose $F(q)$ is decomposed as stated in the Theorem 3.8. Then, there exists constant unitary matrices U&VU\ \&\ V and a regular function $g:\Omega\rightarrow\mathbb{H}$ such that
$F(q)=U\begin{bmatrix}\|F(q_{0})\|&0\\ 0&g(q)\end{bmatrix}V^{*}=UA(q)V^{*}.$
Since $U$ and $V$ are unitary matrices $\|F(q)\|=\|A(q)\|=\max\{\|F(q_{0})\|,\|g(q)\|\}$. If $\|F(q_{0})\|>\|g(q)\|$, then $\|F(q)\|=\|A(q)\|=\|F(q_{0})\|$. This is not possible as $\|F(q)\|$ is non-constant. On the otherhand, if $\|g(q)\|>\|F(q_{0})\|$, then again we have $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\Omega$, which in turn implies that $\|F(q)\|<\|g(q)\|=\|F(q)\|$ which is again a contradiction. Thus, a decomposition as stated in Theorem 3.8 does not hold for $F(q)$, thereby elucidating that the point $q_{0}$ must be inside the domain $\Omega$ in Theorem 3.8.
Remark 3.11.
It should be pointed out that we have assumed that the point $q_{0}\in\Omega_{I}$ (the point where the maximum is attained for $|F(q)|$) some $I\in\mathbb{S}$ in the theorems proved in this section and the previous one. However, this is merely a technical issue and the proofs will go through verbatim, with appropriate modifications, if $q_{0}\in\Omega$. This is because for any $q\in\mathbb{H}$, we can choose $I,J\in\mathbb{S}$ such that $q\in L_{I}$ by taking $v=x_{1}i+x_{2}j+x_{3}k$ (the imaginary part of $q$) and by letting $I=\dfrac{v}{|v|}$ so that $q=x_{0}+|v|I\in L_{I}$.
3.3. Maximum principles for singular values
It is a known fact that singular values and the singular value decomposition of complex matrices have numerous applications. Singular values of a quaternion matrix $A$ are defined to be the positive square roots of the eigenvalues of $A^{*}A$, where $A^{*}$ denotes the conjugate transpose of $A$. It is also not hard to verify that the singular values of the complex adjoint matrix $\widetilde{A}$ are repeated twice, and hence the singular values of $A$ are taken to be the same as it appears for the complex adjoint matrix, without repetition. The SVD of quaternion matrices also finds applications in areas such as image processing (see for instance [11]; also refer to [12] for details on SVD of quaternion matrices). This helps us to prove yet another maximum principle for a matrix-valued regular function $F$, when the singular values of $F(q)$ attains its maximum in $\Omega_{I}$.
Theorem 3.12.
Let $\Omega$ be a slice domain in $\mathbb{H}$ and $F:\Omega\rightarrow M_{n}(\mathbb{H})$ a regular matrix-valued function. If for each $k=1,2,\cdots,n$, the function $q\rightarrow s_{k}(F(q))$ attains its maximum value on $\Omega_{I}$, then $F(q)$ is a constant on $\Omega$.
Proof.
Let
$F(q)=\begin{bmatrix}f_{11}&\cdots&f_{1n}\\ \vdots&&\vdots\\ f_{n1}&\cdots&f_{nn}\end{bmatrix}$
By the splitting lemma (Lemma 2.3), we can write the $(ij)^{th}$ entry of $F(q)$ for some $I\in\mathbb{S}$ as follows: $f_{I}(ij)(q)=g_{ij}(q)+h_{ij}(q)J$ where $g_{ij}$ and $h_{ij}$ are holomorphic functions from $\Omega_{I}$ to $L_{I}$, with $J\perp I$. Thus,
$F_{I}(q)=\begin{bmatrix}f_{11}&\cdots&f_{1n}\\ \vdots&&\vdots\\ f_{n1}&\cdots&f_{nn}\end{bmatrix}=\begin{bmatrix}g_{11}&\cdots&g_{1n}\\ \vdots&&\vdots\\ g_{n1}&\cdots&g_{nn}\end{bmatrix}+\begin{bmatrix}h_{11}&\cdots&h_{1n}\\ \vdots&&\vdots\\ h_{n1}&\cdots&h_{nn}\end{bmatrix}J=G(q)+H(q)J.$
The complex adjoint matrix of $F_{I}(q)$ is the $(2n)\times(2n)$ matrix $\widetilde{F}(q)$ given by
$\widetilde{F}(z)=\begin{bmatrix}[r]G&H\\ -\overline{H}&\overline{G}\end{bmatrix}$
for all $z\in\Omega_{I}$. By the observation stated preceding this theorem, the singular values of $\widetilde{F}(q)$ is repeated twice as that of the singular values of $F(q)$; therefore for each $k=1,2,\cdots,n$, if $q\rightarrow s_{k}(F(q))$ attains its maximum in $\Omega_{I}$, then for each $k=1,2,\cdots,2n,\ z\rightarrow s_{k}(\widetilde{F}(z))$ attains its maximum in $\Omega_{I}$. Then, by the complex version of singular value maximum principle (Theorem $5$ of [7]), we see that $\widetilde{F}(z)$ is a constant on $\Omega_{I}$. This means the functions $G$ and $H$ are constant on $\Omega_{I}$. I n other words, $F_{I}$ is a constant on $\Omega_{I}$. The desired conclusion then follows from the identity principle (Theorem 2.6). ∎
As the following example illustrates, regularity of the entries of the matrix-valued function cannot be dispensed with in the above theorem.
Example 3.13.
Let $F:\mathbb{B}\rightarrow M_{n}(\mathbb{H})$ be defined by $F(q)=\begin{bmatrix}x_{0}+x_{2}j&0\\ 0&1\end{bmatrix}$. Notice that $F$ is not a regular function as the entry $x_{0}+x_{2}j$ is not regular. The singular values of $F$ can be easily evaluated to be $\sqrt{x_{0}^{2}+x_{2}^{2}}$ and $1$ respectively. In this case, the function $s:\mathbb{B}\rightarrow\mathbb{H}$ defined by $s(q)$ (the singular value of $F(q)$) has maximum value as $1$ for any $q\in\mathbb{H}$. However, the function $F(q)$ is a non-constant function.
We end this section by pointing out that the singular values of a matrix-valued function cannot attain its minimum value at a point in the domain unless the matrix-valued function is not invertible at that point.
Theorem 3.14.
Let $F:\Omega\rightarrow M_{n}(\mathbb{H})$ be a non-constant regular function on a symmetric slice domain $\Omega$ in $\mathbb{H}$. Suppose there exist $q_{0}\in\Omega_{I}$ such that for all $k=1,\cdots,n$, the singular values $s_{k}(F(q))$ achieves its minimum at $q_{0}$. Then $F(q_{0})$ is not invertible.
Proof.
Since the singular values $s_{k}(\widetilde{F}(q))$ of the complex adjoint matrix $\widetilde{F}(q)$ are repeated twice as that of the singular values of $F(q)$, if $s_{k}(F(q))$ attains its minimum at $q_{0}\in\Omega_{I}$ for all $1\leq k\leq n$, then $s_{k}(\widetilde{F}(q))$ attains its minimum at $q_{0}$. It then follows from Theorem $9$ of [7] that $\widetilde{F}(q_{0})$ is not invertible, which in turn implies that $F(q_{0})$ is not invertible. ∎
3.4. Approximation of a quaternionic matrix-valued function
This is the last section of the paper. We prove as an application of Theorem 3.8, a Caratheodory-Rudin type approximation theorem. We begin with the following two definitions.
Definition 3.15.
Let $\Omega$ be subset of $\mathbb{H}$. A function $f:\Omega\rightarrow\mathbb{H}$ is said to be a rational function if $f(q)=h(q)^{-*}*g(q)$ where $h(q)$ and $g(q)$ are (left) polynomials with coefficients from quaternions such that $h(q)$ has no zeros in $\Omega$.
Definition 3.16.
A function $f:\mathbb{B}\rightarrow M_{n}(\mathbb{H})$ is called an inner function if the boundary values of the function on the unit ball are unitary matrices everywhere.
We are now in a position to prove a Caratheodory-Rudin type approximation result. The proof uses a Runge type approximation theorem over quaternions. Runge theorem for slice hyperholomorphic functions was recently studied in [5].
Theorem 3.17.
Let $\mathbb{B}$ be the open unit ball in $\mathbb{H}$ and $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ be a regular function on $\mathbb{B}$. Then $f$ can be approximated by rational inner functions.
Proof.
By the splitting lemma (Lemma 2.3), any regular function $f$ can be written as $f_{J}=g+hK$, where $J,K\in\mathbb{S},\ J\perp K$ and $g,h:\mathbb{B}_{J}\rightarrow\overline{\mathbb{B}_{J}}$ being holomorphic functions on $\mathbb{B}_{J}$. By the Caratheodory-Rudin theorem for the complex case (see [2] for details), $g$ and $h$ can be approximated by rational inner functions in field of complex numbers, say $r_{1}$ and $r_{2}$, such that for $\epsilon>0$, we have $|g(z)-r_{1}(z)|<\epsilon/4$ and $|h(z)-r_{2}(z)|<\epsilon/4$ for all $z=x+yJ\in\mathbb{B}_{J}$. We can extend $r(x+yJ):=r_{1}(x+yJ)+r_{2}(x+yJ)K$ to the entire unit ball $\mathbb{B}$ by extension, say, $r(z)$ for $I\in\mathbb{S}$ (see Proposition $1.3.2$ of [8] for details) for $I\in\mathbb{S}$. Then,
$ext(r(q))=r(x+yI)=\frac{1}{2}\left[r(x+yJ)+r(x-yJ)+IJ\left[r(x-yJ)-r(x+yJ)\right]\right].$
Consider
$|f(q)-r(q)|$ $=\Bigg|\frac{1}{2}[f(x+yJ)+f(x-yJ)]+IJ\left[f(x-yJ)-f(x+yJ)\right]$
$-[r(x+yJ)+r(x-yJ)+IJ[r(x-yJ)-r(x+yJ)]]\Bigg|.$
Once again by the splitting lemma, the previous equation can be expressed as
$|f(q)-r(q)|$ $=\Bigg|\frac{1}{2}\Bigg[(1-IJ)\Big[g(x+yJ)-r_{1}(x+yJ)+(h(x+yJ)-r_{2}(x+yJ))K\Big]$
$+(1+IJ)\Big[g(x-yJ)-r_{1}(x-yJ)+(h(x-yJ)-r_{2}(x-yJ))K\Big]\Bigg]\Bigg|$
$\leq\Big|g(x+yJ)-r_{1}(x+yJ)\Big|+\Big|h(x+yJ)-r_{2}(x+yJ)\Big|$
$+\Big|g(x-yJ)-r_{1}(x-yJ)\Big|+\Big|h(x-yJ)-r_{2}(x-yJ)\Big|$
$\leq\epsilon.$
We thus conclude that $f$ can be approximated by rational inner functions over $\mathbb{H}$, as desired. ∎
Remark 3.18.
As defined in [3], regular functions that satisfy the hypothesis of the above theorem are said to be in quaternionic Schur class (see Definition $1.1$ of [3]). We have thus proved that any function that belongs to the quaternionic Schur class can be approximated by rational inner functions over $\mathbb{H}$. As pointed out in [3], both right and left Schur classes over quaternions coincide. As a final remark, we wish to point out that the quaternionic Schur class was studied in detail by Alpay et al (see Theorem $3.3$ and its consequences of [1]) and Bolotnikov (see Section $4$ of [4]).
We end this paper by proving the following result using the maximum norm principle.
Theorem 3.19.
Let $F:\mathbb{B}\rightarrow M_{2}(\mathbb{H})$ be a regular function such that $\|F(q)\|=1$. Then $F$ can be approximated by $M_{2}(\mathbb{H})$-valued rational inner functions.
Proof.
Since $\|F(q)\|=1$, Theorem 3.8 tells us that there exists $2\times 2$ constant unitary matrices $U$ and $V$ and a regular function $G:\mathbb{B}\rightarrow M_{1}(\mathbb{H})\cong\mathbb{H}$ such that
$F(q)=U\begin{bmatrix}1&0\\ 0&G(z)\end{bmatrix}V^{*}.$
The Theorem 3.17 will then imply that there is an approximation of $G$ by rational inner functions, and hence $F$ can also be approximated by matrix-valued rational inner functions. ∎
Declarations:
  • Both Sachindranath Jayaraman and Dhashna T. Pillai have equal contribution in this work and declare that there is no conflict of interest.
  • No funding was received for conducting this study.
  • The authors also declare that there is no data used in this research.